OpenStudy (kidthatbro8):

Solve for x: −3|2x + 6| = −12 x = 1 and x = 5 x = −1 and x = −5 x = −9 and x = 3 No solution could someone explain to me how to do this so that i don't have to come back ?

OpenStudy (anonymous):

idk sorry

OpenStudy (kidthatbro8):

really?,,,,,,,well, i really need help so.....

OpenStudy (anonymous):

seriously. sorry lmaoo

jimthompson5910 (jim_thompson5910):

first divide both sides by -3 −3|2x + 6| = −12 [−3|2x + 6|]/(-3) = −12/(-3) |2x+6| = 4

jimthompson5910 (jim_thompson5910):

Then break up |2x+6| = 4 into 2x+6 = 4 or 2x+6 = -4 solve each equation for x to get your two answers

OpenStudy (kidthatbro8):

oh, thank you ! at least i know how to solve it.